OpenStudy (anonymous):

The length of a rectangle is 4 cm more than twice the width. the perimeter is 80 cm. Find the dimensions of the rectangle

OpenStudy (ajprincess):

Perimeter=2(l+w) where l-length and w-width l=w+4 P=2(w+4+w) 80=2(2w+4) 80/2=2w+4 40=2w+4 solve for w. Then substitute the value of w in l=w+4 to find l. Can u do?

OpenStudy (anonymous):

So L equals 22, w= 18 and the area would be 396?

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

okay how about -1/4(-36+12x)=?

OpenStudy (anonymous):

one fourth btw, not -1 divided by 4...

OpenStudy (ajprincess):

you mean ur question is 1/4(-36+12x)=?

OpenStudy (anonymous):

yeaaa!

OpenStudy (jiteshmeghwal9):

first of all use distributive property to solve denominator first:)

OpenStudy (anonymous):

okay. so what would that be?

OpenStudy (jiteshmeghwal9):

\[\frac{1}{-144+48x}\]

OpenStudy (anonymous):

huh? im sooo confused.

OpenStudy (anonymous):

ohhh nevermind. so you did -4 and distributed it to -36 and +12x

OpenStudy (jiteshmeghwal9):

don't u know distributive property ???????

OpenStudy (anonymous):

yessss

OpenStudy (jiteshmeghwal9):

so i use distributive property to solve denominator means i distribute 4 btw -36+12x

OpenStudy (jiteshmeghwal9):

gt it @maherman97 ??????

OpenStudy (anonymous):

yeahh i know i figured that out. so then what?

OpenStudy (jiteshmeghwal9):

so now it must be ur answer becoz it can't be solved further:)

OpenStudy (anonymous):

or is that simplest form? ohh okay thanks man

OpenStudy (jiteshmeghwal9):

yw:) Anytime